[figure: 9.7 cm]
NEXAFS Study of Pr_{0.65}Ca_{0.35}MnO_{3} Films
Y. S. Lee^{a}, V. G. Prokhorov^{b}, H. J. Shin^{c}, J. M. Kim^{d},
H. G. Park^{e}, and Y. P. Lee^{e}

aDiv. of Inf. Comm. & Comp. Eng., Hanbat Nat’l Univ. Daejeon 305-719, Korea

bInstitute of Metal Physics, NASu, Kiev, 03142, Ukraine c Pohang Accelerator Laboratory, Pohang Kyungbuk 790-784, Korea d Department of Physics, Hoseo University, Asan Chungnam 336-795, Korea e Department of Physics, Hanyang University, Seoul 133-791, Korea

The magnetic and electronic properties of the colossal magnetoresistnace (CMR) compounds have been traditionally explained by the double-exchange model which considers the transfer of an electron (or a hole) between neighboring Mn^{3+} and Mn^{4+} ions. Pr_{0.65}Ca_{0.35}MnO_{3} compound belongs to a family of perovskite-like manganite system which shows the CMR properties. The Pr_{0.65}Ca_{0.35}MnO_{3} system consists formally of the Mn^{3+} and Mn^{4+} ions. The magnetic and transport proterties of perovskite manganite system can be varied by changing the Mn^{3+}/Mn^{4+} ratio. The prepared Pr_{0.65}Ca_{0.35}MnO_{3} films have different structures according to the substrate structure through the chemical compositions are same for all the films, which are also induced by the change in the Mn^{3+}/Mn^{4+} ratio. In this study, the polycrystalline Pr_{0.65}Ca_{0.35}MnO_{3} films were prepared by laser ablation on various substrates, i.e., Polico, YSZ, SiTiN, and LaAlO_{3}(100) single crystal. We investigated the properties according to the substrates by using Pr M_{4,5}-, Ca L_{2,3}-, Mn L_{2,3}-, and O K -edge x-ray absorption spectroscopy.

[figure: 13.4 cm]

Subject category :

2. Magnetic Films, Surfaces and Multilayers

Presentation mode : poster

Corresponding author : Y. S. Lee

Address for correspondence : Div. of Information Communication & Computer Eng., Hanbat Nat’l Univ. Daejeon 305-719, Korea

Email address : yslee@hanbat.ac.kr